OpenStudy (anonymous):

The sum of a two digit numer is ten. If the digits are reversed, the new umber is fifty-four moe than the original number. Find the original number.

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

let digits of original number be xy in that order

OpenStudy (anonymous):

so x+y=10 ( one eqn )

OpenStudy (anonymous):

reversing the digits ( yx ) gives 54 more than the original

OpenStudy (anonymous):

ie (y * 10 ) + x = 54 + ( x*10) +y

OpenStudy (anonymous):

two eqns , two unknowns

OpenStudy (anonymous):

you can solve

OpenStudy (anonymous):

remember the number "xy" = (x*10) + y because we have x in the 10'splace and y in the ones place

OpenStudy (anonymous):

thats a slightly different question I havent seen before

OpenStudy (anonymous):

so x+y=10 and 9x -9y +54 =0 solve those simultaneously to find x and y , you can do that

OpenStudy (anonymous):

then the original number will be x followed by y so if you find x=4 and y=6 the orginal; number would have been 46

OpenStudy (anonymous):

substitutoin right

OpenStudy (anonymous):

could do anything you want, doesnt really matter

OpenStudy (anonymous):

you could divide the 2 eqn by 9 , and then add the eqns and solve by elimination